Write an equation of the line containing the given point and parallel of the given line. Express your answer in the form y=mx+b
(5,-7); 7x+2y=3

Answer by mananth(16949)

(5,-7);
7x+2y=3
-7x
2y=-7x+3
/2
y=(-7/2)x+3/2
...
compare with y= mx+b
m=(-7/2)
equation of line parallel to the above line will have same slope.
Point(-5,7) slope = (-7/2)
Plug these values to get b
y = mx +b
7=(-7/2)*(-5)+b
7=35/2+b
7-35/2=b
(14-35)/2 = b
-21/2 = b
...
Equaqtion will be
y = (-7/2)x-21/2
...

...
m.ananth@hotmail.ca